How long is a guy wire that reaches from the top of a 15-ft pole to a point on the ground 10 feet from the bottom of the pole? Find the length to the nearest tenth of a foot.
Answer by texttutoring(324)

Use Pythagorean Theorem: a^2 + b^2 = c^2
Draw a triangle, with the vertical 15 ft pole and the horizontal 10 ft line on the ground making up the legs of the triangle.
The hypotenuse of the triangle would be the wire.

c^2 = 15^2 + 10^2
c^2 = 225 + 100
c^2 = 325
c = 18.0 ft

The wire is 18.0 feet long.